OpenStudy (anonymous):

find the fourth term of the geometric sequence whose first term is 4 and whose common ratio is -2

OpenStudy (anonymous):

-32

OpenStudy (anonymous):

4*(-2)^3 = -8*4=-32

OpenStudy (anonymous):

Lia it is -32

OpenStudy (anonymous):

asdflkjsadfd. i give up.

OpenStudy (anonymous):

wt?

OpenStudy (anonymous):

nothing. i apparently can't do simple math. it's kind of pathetic.

OpenStudy (anonymous):

\[ar^n\]